# Revealing Superconducting Chiral Edge Modes via Resistance Distributions

### v1

@@ Abstract
For the junction configuration of Fig. 1(d), the results are presented in Fig. 3(a,b). When $L\gtrsim 2\ell_{d}$, both the mean and fluctuations of $T$ remain constant as $L$ varies, with $\overline{T}$ aligning with its value in the clean limit. Furthermore, the distribution of $T$ follows a normal distribution, consistent with Case (iii), confirming that this distribution arises from the random coupling of Majorana modes near the junction corners. Next, the results for the QAH-SC junction of Fig. 1(f) are shown in Fig. 3(c,d). While $\overline{T}$ exhibits oscillations with decaying amplitude—similar to the $N=2$ case—it does not necessarily converge to zero, depending on the incident electron energy $\varepsilon$. Furthermore, for sufficiently large $L$, such that $L\gtrsim 10^{4}$, the distribution transitions to a generalized arcsine form, as predicted in Case (iv). The left peak is broadened due to the neglected fluctuations of the rotation axis in the analytical analysis, which introduce slight variations in the lower bound of the distribution.
@@
Decoherence. Two additional factors, loss and decoherence, play a significant role in shaping experimental observations beyond the effects of disorder [26]. Loss occurs when particles tunnel into magnetic vortices and escape through the vortex line to the substrate [48, 46, 74]. This process reduces the norm of $\Psi_{f}$, thus compressing the distribution of $T$ without altering its overall shape. Decoherence, on the other hand, results from inelastic scattering with other gapless degrees of freedom, which can be induced by magnetic vortices, ineffective SC proximity, or thermal fluctuations [49]. When the decoherence effect is weak, the effect of decoherence can be conceptualized as the edge mode that has a probability $p$ of undergoing a single inelastic scattering event during its propagation. The resulting effective charge transmission fraction for different scenarios in Fig. 1(a,d,f) is shown in Fig. 4, where the distribution for $N=2$ CBEM remains distinct from that for $N=1$ CBEM, indicating that the characteristic behavior of the distribution in the coherent process persists under sufficiently weak decoherence. As the decoherence effect intensifies, characterized by an increase in $p$ and the occurrence of multiple scattering events, the distribution for the scenario in Fig. 1(a) will become logarithmic divergent peaked at $T=0$, while the scenarios in Fig. 1(d,f) become bell-shaped peaked at finite $T$ [62].
@@
Figure 4: (a,b,c) Decoherence effect in scenarios Fig. 1(a,d,f), respectively. The distribution is calculated from Monte Carlo sampling [62], and we set $p=0.4$. We set the distribution of $T$ without decoherence follows $U[-1,1]$, $\mathcal{N}(0.3,0.05^{2})$ and arcsin distribution with $A=0.6$ for Fig. 1(a,d,f), respectively.
Discussions. The experimentally measurable resistances are determined by $T$. Taking QAH-SC junction shown in Fig. 1(f) as an example. The current is from lead 1 to the grounded SC, namely $I_{1}=-I_{s}=I$, and voltage leads are on electrodes 2, 3, 4, and 5. The downstream resistance [19, 25, 28] is given by the generalized Landauer-Büttiker formula [75]
$R_{d}\equiv R_{3s}=\frac{V_{3}-V_{s}}{I}=\frac{h}{e^{2}}\frac{T}{1-T},$ (6)
where $h$ is the Planck constant, $e$ is the electrical charge. Thus, $T$ can be directly extracted from the measured $R_{d}$. Other nonlocal resistances in different scenarios are provided in [62].
@@
Our theory shows that the probability distribution of charge transmission $T$ provides a signature distinguish between the different topological phases of the SC region. As summarized in Table 1, for sufficiently long junctions, the distributions for $N=1$ and $N=2$ CBEM exhibit distinct characteristics. Moreover, these differences remain robust against weak decoherence. Given the stark differences between these distributions, they provide stronger evidence for phase differentiation, offering extra insight than mean alone, which may be insufficient to distinguish between topological phases of SC region [28].
@@ Acknowledgements.
Acknowledgment. We thank Yuanbo Zhang for enlightening discussions. This work is supported by the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01.

### v2

@@ IV Numerical simulations
For the junction configuration of Fig. 1(d), the results are presented in Figs. 3(a) and 3(b). When $L\gtrsim 2\ell_{d}$, both the mean and fluctuations of $T$ remain constant as $L$ varies, with $\overline{T}$ aligning with its value in the clean limit. Furthermore, the distribution of $T$ follows a normal distribution, consistent with case (iii), confirming that this distribution arises from the random coupling of Majorana modes near the junction corners. The behavior of $R$, the charge transmission probability from lead 3 to lead 5, exhibits a similar behavior [64]. Next, the results for the QAH-SC junction of Fig. 1(f) are shown in Figs. 3(c) and 3(d). While $\overline{T}$ exhibits oscillations with decaying amplitude—similar to the $N=2$ case—it does not necessarily converge to zero, depending on the incident electron energy $\varepsilon$. Furthermore, for sufficiently large $L$, in this parameter setup $L\gtrsim 2\times 10^{4}$, the distribution transitions to a generalized arcsine form, as predicted in case (iv). The left peak is broadened due to the neglected fluctuations of the rotation axis in the analytical analysis, which introduce slight variations in the lower bound of the distribution.
V Effect of decoherence and particle loss
Two additional factors, decoherence and particle loss, play a significant role in shaping experimental observations beyond the effects of disorder [26]. Decoherence refers to the loss of quantum phase coherence of the state $\Psi(x)$ during propagation, arising from two sources: (1) inelastic scattering with other gapless degrees of freedom, which can be induced by magnetic vortices, incomplete SC proximity, or thermal fluctuations [50]; and (2) temperature-induced dephasing. Loss occurs when particles tunnel into magnetic vortices and escape through the vortex line to the substrate [49, 47, 78]. In the following, we discuss the effects of decoherence and loss separately.
V.1 Decoherence effect
We first consider decoherence induced by inelastic scattering. Phenomenologically, this effect can be incorporated by introducing floating leads [79] at the boundaries of the TSC region, which serve as decoherence sources. In the simplest case, with only one inelastic scattering center along the edge, we introduce a parameter $p$ to describe the probability that edge modes enter a floating lead. For a QAH-SC-QAH junction or a QAH-SC junction with $N=2$ CBEMs, the effective charge transmission fraction from lead 2 to lead 3 is given by the generalized Landauer-Büttiker formula [64]:
$T^{\text{eff}}=(1-p)T^{(l,r)}+pT^{(l,f)}T^{(f,r)},$ (23)
where $T^{(l,r)}$ denotes the coherent charge transmission fraction from lead 2 to lead 3, $T^{(l,f)}$ is the fraction from lead 2 to the floating lead $f$, and $T^{(f,r)}$ is the fraction from lead $f$ to lead 3. For a QAH-SC junction with $N=1$ CBEM, the situation is slightly different because the two Majorana modes propagate along distinct edges of the TSC. For simplicity, we consider one floating lead on each edge, assuming the same decoherence probability $p$ for both. The effective charge transmission fraction is then given by
$T^{\text{eff}}=p\left[T^{(l,u)}T^{(u,r)}+T^{(l,d)}T^{(d,r)}\right]+(1-p)T^{(l,r)},$ (24)
where $T^{(l,u)}$ and $T^{(l,d)}$ are the transmission fractions from lead 2 to the upper and lower floating leads, respectively, $T^{(u,r)}$ and $T^{(d,r)}$ are the corresponding fractions from the floating leads to lead 3, and $T^{(l,r)}$ is the coherent transmission fraction from lead 2 directly to lead 3.
@@ V.1 Decoherence effect
When the decoherence effect is weak, the resulting effective charge transmission fraction for different scenarios in Figs. 1(a), 1(d) and 1(f) is shown in Fig. 4, where we focus on the long junction case, i.e., each charge transmission fraction converges to the form of its large $L$ limit. The distribution for $N=2$ CBEMs remains distinct from that for $N=1$ CBEM, indicating that the characteristic behavior of the distribution in the coherent process persists under sufficiently weak decoherence. As the decoherence effect intensifies, characterized by an increase in $p$ and the occurrence of multiple scattering events, the distribution of the scenario in Fig. 1(a) will become logarithmic-divergent peaked at $T=0$, while the scenarios in Fig. 1(d) and 1(f) become bell-shape peaked at finite $T$ [64].
@@
Figure 4: (a), (b), (c) Decoherence effect in scenarios Figs. 1(a), 1(d) and 1(f), respectively. The distribution is calculated from Monte Carlo sampling [64], and we set $p=0.4$. We set the distribution of $T$ without decoherence follows U[1,1]U[-1,1], $\mathcal{N}(0.3,0.05^{2})$ and arcsin distribution with $A=0.6$ for Figs. 1(a), 1(d) and 1(f), respectively.
If multiple inelastic scattering centers exist, additional floating leads need to be introduced, which effectively divides the TSC region into several segments and leads to a more complicated multi-terminal formula. We consider the simple case when $p=1$, which means that edge modes fully enter every floating lead and completely lose coherence. Then Eq. (23) reduces to [50]
$T^{\text{eff}}=T_{l,1}T_{1,2}\cdots T_{m,r},$ (25)
when there are $m$ leads at the edge, where the subscripts $1$ to $m$ represent virtual floating leads. For QAH-SC junction with $N=1$ CBEM, the effective transmission fraction is instead given by the sum of two paths:
$T^{\text{eff}}=T^{u}_{l,1}T^{u}_{1,2}\cdots T^{u}_{m,r}+T^{d}_{l,1}T^{d}_{1,2}\cdots T^{d}_{m,r}.$ (26)
When more floating leads are introduced, the distribution of $T^{\text{eff}}$ will become increasingly sharply peaked, and its peak position will shift toward zero [64].
We now briefly address the decoherence induced by temperature. Unlike the discrete nature of inelastic scattering centers, the temperature effect can be modeled as a continuous decoherence source by attaching floating leads to every site along the edge while keeping the parameter $p$ very small. Intuitively, in the Bloch-sphere picture, decoherence transforms a pure state located on the Bloch surface into a mixed state inside the sphere. In the case of temperature-induced decoherence, this corresponds to the point gradually shrinking toward the center of the Bloch sphere as the edge state propagates. Consequently, the overall effect is a narrowing of the distribution of $T$, without changing its qualitative form. Since our objective is to use the distribution to distinguish different edge scenarios, temperature-induced decoherence does not qualitatively affect the results, provided the effect remains weak, i.e., low temperature.
V.2 Particle loss
The phenomenological treatment of particle loss is straightforward. Since this process corresponds to particles that escape from the device, the parameters $q$ (with values between 0 and 1) can be introduced to describe the reduction of particle number: $\Psi_{f}=q_{m}\cdots q_{2}q_{1}\Psi^{0}_{f}$, where $m$ is the number of vortices and $\Psi^{0}_{f}$ denotes the state in the absence of particle loss. This factor merely suppresses the amplitude of $T$, which means that particle loss compresses the distribution of $T$ without altering its overall form, similar to the effect of temperature-induced decoherence. Therefore, only inelastic scattering modifies the functional form of the distribution and exerts a qualitative influence, whereas temperature and particle loss have only quantitative effects.
VI Discussion and conclusion
Finally, we discuss the relation between experimentally measurable quantities and the charge transmission fraction introduced above. We restrict the discussion to the linear-response regime, corresponding to a small applied bias. Within this approximation, the charge transmission function can be defined as [80]
$\mathcal{T}(\varepsilon)=\int T(\xi)\left[-\frac{\partial f_{\text{FD}}(\xi-\varepsilon)}{\partial\xi}\right]\mathrm{d}\xi,$ (27)
where $f_{\text{FD}}$ is the Fermi-Dirac distribution, $\varepsilon$ is the incident energy, and we have explicitly written the energy dependence of $T$ defined in previous sections. Roughly the kernel $\frac{\partial f_{\text{FD}}(\xi-\varepsilon)}{\partial\xi}=\left(4k_{B}\tau\cosh^{2}\frac{\xi-\varepsilon}{2k_{B}\tau}\right)^{-1}$ effectively averages $T(\xi)$ over an energy window of order $\varepsilon\pm k_{B}\tau$, where $\tau$ denotes the temperature. Taking the QAH-SC junction shown in Fig. 1(f) as an example, the current is from lead 1 to the grounded SC, namely $I_{1}=-I_{s}=I$, and voltage leads are on electrodes 2, 3, 4, and 5. Under linear-response approximation, the experimentally measurable downstream resistance [19, 25, 28], a nonlocal differential resistance, is given by the generalized Landauer-Büttiker formula [81]
$R_{d}\equiv R_{3s}=\frac{V_{3}-V_{s}}{I}=\frac{h}{e^{2}}\frac{\mathcal{T}}{1-\mathcal{T}},$ (28)
where $h$ is the Planck constant and $e$ is the electrical charge. Thus $\mathcal{T}$ can be directly extracted from the measured $R_{d}$. Other nonlocal resistances in different scenarios are provided in Ref. [64]. In general, the distribution of $\mathcal{T}$ can differ from that of $T$; the mapping between them—given by the energy convolution with the kernel $\partial f_{\text{FD}}/\partial\xi$—is analytically complicated and depends on the detailed energy dependence of $T(\xi)$ and on the temperature $\tau$. However, if the smearing window $k_{B}\tau$ is much smaller than the superconducting gap, or $T(\xi)$ is not sensitive to $\xi$ (by, e.g., decoherence effects), then $T(\xi)$ varies only weakly and thus the distribution of $\mathcal{T}(\varepsilon)$ approaches that of $T(\varepsilon)$. Under this situation, thermal smearing effect is negligible; therefore the distinct behaviors of $\mathcal{T}$ for the $N=1$ and $N=2$ cases remain observable.
@@ VI Discussion and conclusion
In conclusion, our theory shows that the probability distribution of charge transmission $T$ provides a signature distinguished between the different topological phases of the SC region. As summarized in Table 1, for sufficiently long junctions, the distributions for $N=1$ and $N=2$ CBEMs exhibit distinct characteristics. Moreover, these differences remain robust against weak decoherence. Given the stark differences between these distributions, they provide stronger evidence for phase differentiation, offering extra insight than mean alone, which may be insufficient to distinguish between topological phases of the SC region [28].
@@ Acknowledgements.
We thank Yuanbo Zhang for enlightening discussions. This work is supported by the National Key Research Program of China under Grant No. 2025YFA1411400, the Natural Science Foundation of China through Grants No. 12350404 and No. 12174066, the Innovation Program for Quantum Science and Technology through Grant No. 2021ZD0302600, the Science and Technology Commission of Shanghai Municipality under Grants No. 23JC1400600, No. 24LZ1400100 and No. 2019SHZDZX01, and is sponsored by the “Shuguang Program” supported by the Shanghai Education Development Foundation and Shanghai Municipal Education Commission.